Please help with this problem.

A first year Calculus class test gave the following information: Mean of 63 and a Standard Deviation of 12. Answer the questions:
A. If 800 took the test how many passes with a 51% or higher?
b. What is the standard score of 39%?
C. If a student got and 82%, what is the standard score?

Answer by nabla(475)

a. 63-51=12 (1 standard deviation below the mean). This is asking for 50%+34%=84% of the tests. .84*800=672. In words, it was asking for the amount of tests scoring higher than 1 standard deviation below the mean.
b. z=(39-63)/12=-2. The score lies precisely 2 standard deviations below the mean.
c. z=(82-63)/12=1.58333...